Construct all random samples consisting three observations from the given data. Arrange the observations in ascending order without replacement and repetition.


86 89 92 95 98


1
Expert's answer
2021-06-14T15:11:07-0400

N=5

n=3

"Number\; of \; samples = \frac{5!}{3!(5-3)!} = \frac{20}{2}=10"

Samples:

86 89 92

86 89 95

86 89 98

86 92 95

86 92 98

86 95 98

89 92 95

89 92 98

89 95 98

92 95 98